OpenStudy (anonymous):

the dress store is having a sale where all merchandise is 1/4 off. a woman buys $54 merchandise at the sale price. what would she have had to pay at the regular price ?

OpenStudy (anonymous):

?

OpenStudy (shinalcantara):

knowing that $54 is already the discounted rate, then you can say that it is 3/4 of the original price. getting the original price you'll have the equation as: let 'x' be the original price \[54 = \frac{ 3 }{ 4 }x\] find 'x'

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

hello?

OpenStudy (shinalcantara):

cross multiply 4 to the left side and you'll have: 54(4) = 3x or 3x = 54(4) 3x = 216 divide both sides of the equation with 3 x = 216/3 x = 72